OpenStudy (anonymous):

Explain how to write the equation y – 2 = 5(x + 7) in slope-intercept form.

OpenStudy (anonymous):

You must carry out the order of operations and solve for Y.

OpenStudy (anonymous):

1. Distribute the 5 to what is in the parenthesis Ex. y – 2 = 5x + 35 2. Since the 2 is negative you add it to both sides of the equal sign Ex. y – 2 = 5x + 35 +2 +2 3. y = 5x + 37